If a force directed east acts on stationary 5.0kg mass , what are its acceleration displacement , and velocity after 10s ?
1
Expert's answer
2020-12-06T17:29:33-0500
"a=\frac{F}{m}=\frac{15}{5}=3\frac{m}{s^2}\\v=at=(3)(10)=30\frac{m}{s}
\\s=0.5at^2=0.5(3)(10)^2=150\ m"